OpenStudy (broskishelleh):

Task 2—Non-Linear Systems of Equations Create a system of equations that includes one linear equation and one quadratic equation. Part 1. Show all work to solving your system of equations algebraically. Part 2. Graph your system of equations, and show the solution graphically to verify your solution

OpenStudy (broskishelleh):

@dan815 @ganeshie8 @iambatman

OpenStudy (anonymous):

Make it very easy, just use something you can think of in your head. Easiest quadratic equation, x^2. Easiest linear equation, x. They're also easy to graph and to find their solution, set x^2 = x. Divide each side by x, x = 1. When doing systems of linear and nonlinear equations, just think what the easiest example to satisfy the problem is!

OpenStudy (solomonzelman):

Basically a parabola and a line.

OpenStudy (broskishelleh):

I needed an f(x) = equation, and a y = mx + b forms

OpenStudy (anonymous):

Just do what solomon said

OpenStudy (broskishelleh):

? I don't know how to make a parabola

OpenStudy (solomonzelman):

Make anything up! Any line and any parabola (opening up or down)

OpenStudy (anonymous):

f(x) = x^2

OpenStudy (anonymous):

y = x

OpenStudy (anonymous):

Here, I"ll pick 2 that I know work. f(x) = 2x + 1 y = x^2 + x - 2

OpenStudy (solomonzelman):

I would try to have b=2,4, or other EVEN number.

OpenStudy (solomonzelman):

I mean in ax^2+bx+c=0

OpenStudy (broskishelleh):

I thank you all! Thank you so much, but can you explain your system @adam.bloedel

OpenStudy (broskishelleh):

Please

OpenStudy (solomonzelman):

I would choose y=2x+2 f(x)=x²+4x+6

OpenStudy (anonymous):

Yeah those are good :)

OpenStudy (broskishelleh):

Well, by explain, I meant prove it was verified

OpenStudy (anonymous):

I would listen to these guys. They know more about how equations act than I do. They can give in-depth nearly professional answers. I know @SolomonZelman and Batman are great at what they do. Listen to Solomon step you through it.

OpenStudy (broskishelleh):

Will do, thank you!

OpenStudy (solomonzelman):

Okay. `y=x+2` (actually 1x, not 2x) would be just a line. It is same as y=x but shifted 2 units up and easy to work with. And `f(x)=x²+4x+6`

OpenStudy (solomonzelman):

`f(x) = x²+4x+6` → `f(x) = x²+4x+4+2` → `f(x) = (x+2)²+2`

OpenStudy (broskishelleh):

They are verified as to be a proper system? f(x) = (x+2)²+2 and y = x + 2

OpenStudy (solomonzelman):

And there are you parabola f(x)=(x+2)²+2 and a line y=x+2

OpenStudy (solomonzelman):

You can check that algebraically.

OpenStudy (broskishelleh):

After this question, can I ask you guys something?

OpenStudy (anonymous):

Do you know how to graph the equations?

OpenStudy (broskishelleh):

Yes

OpenStudy (broskishelleh):

I did

OpenStudy (solomonzelman):

I changed y=x+2 to y=x+6 if you don't mind.

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/mt7rqitrms

OpenStudy (broskishelleh):

How can I show you? Well, it looks like this: The parabola's minimum is -2,has no x-intercept, and y = 6

OpenStudy (broskishelleh):

Oh, we changed it, ok

OpenStudy (broskishelleh):

I have to draw it again lol, thanks for patience

OpenStudy (solomonzelman):

Yes, I drew it, https://www.desmos.com/calculator/mt7rqitrms :)

OpenStudy (broskishelleh):

They intercept!

OpenStudy (broskishelleh):

The line: x = (-6, 0) y = (0, 6) y = 1x + 6 Slope = 1 6 = y

OpenStudy (broskishelleh):

So this is verified I believe

OpenStudy (broskishelleh):

After this question, can I ask you guys something?

OpenStudy (broskishelleh):

It has to do with my math course, you two seem to be extremely smart about math...